OpenStudy (anonymous):

solve for u u^2-5u+6=0

OpenStudy (anonymous):

Simplifying u2 + -5u + -6 = 0 Reorder the terms: -6 + -5u + u2 = 0 Solving -6 + -5u + u2 = 0 Solving for variable 'u'. Factor a trinomial. (-1 + -1u)(6 + -1u) = 0 Subproblem 1 Set the factor '(-1 + -1u)' equal to zero and attempt to solve: Simplifying -1 + -1u = 0 Solving -1 + -1u = 0 Move all terms containing u to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + -1u = 0 + 1 Combine like terms: -1 + 1 = 0 0 + -1u = 0 + 1 -1u = 0 + 1 Combine like terms: 0 + 1 = 1 -1u = 1 Divide each side by '-1'. u = -1 Simplifying u = -1 Subproblem 2 Set the factor '(6 + -1u)' equal to zero and attempt to solve: Simplifying 6 + -1u = 0 Solving 6 + -1u = 0 Move all terms containing u to the left, all other terms to the right. Add '-6' to each side of the equation. 6 + -6 + -1u = 0 + -6 Combine like terms: 6 + -6 = 0 0 + -1u = 0 + -6 -1u = 0 + -6 Combine like terms: 0 + -6 = -6 -1u = -6 Divide each side by '-1'. u = 6 Simplifying u = 6 Solution u = {-1, 6}